OpenStudy (harrisonshark):

What is the distance between the points (2, 5) and (5, 7)?

OpenStudy (anonymous):

d = sqrt[ (x2-x1)^2 + (y2-y1)^2]

OpenStudy (harrisonshark):

the answer are sqrt 10 sqrt 13 and 5

OpenStudy (harrisonshark):

either of them

OpenStudy (harrisonshark):

@douglaswinslowcooper

OpenStudy (anonymous):

d = sqrt(9 + 4) = sqrt(13)

OpenStudy (harrisonshark):

thank ou